# Better Together? A Field Experiment on

### v2

@@ 2. Scored 13,115 7,021 7,345
3. First Scored Referral in Trial 6,852 6,852 4,213
@@ 8. Final Sample, Excluding Observations with Overwritten Status 4,681 4,681 2,832
Full Sample Hospital Outcomes Sample
Female 0.000 0.008
[0.014] [0.017]
Age at Initial Randomization -0.002 -0.003
[0.001] [0.001]
White 0.049 0.010
[0.070] [0.074]
Black -0.048 -0.102
[0.082] [0.088]
Missing Race -0.000 -0.028
[0.077] [0.085]
Hispanic 0.034 0.058
[0.029] [0.033]
Missing Ethnicity 0.013 0.006
[0.037] [0.044]
Algorithm Score -0.003 0.000
[0.002] [0.002]
Mandated Reporter -0.011 -0.014
[0.024] [0.027]
Child Reported in Danger 0.038 0.047
[0.068] [0.078]
Team 3 0.039 0.048
[0.025] [0.030]
Team 2 0.004 0.016
[0.026] [0.030]
Household Size 0.001 0.001
[0.014] [0.016]
Number of Alleged Child Victims 0.000 -0.009
[0.017] [0.019]
Active -0.014 -0.047
[0.034] [0.037]
Number of Previous Referrals 0.001 0.001
[0.002] [0.002]
Harm Index 180 Days Prior to Referral 0.005 0.004
[0.008] [0.010]
Any Prior Hospital Claim -0.037∗∗ -0.052∗∗
[0.018] [0.020]
Number of Prior Hospital Claims 0.002 0.004
[0.002] [0.002]
Hospital Outcomes Sample 0.002
[0.023]
Randomization Controls Yes Yes
Share Randomized to Treatment 0.55 0.55
Observations 4,681 3,431
Randomization Clusters 2,810 2,055
F-statistic 0.965 1.421
p-value of F-test 0.503 0.106
Notes: This table presents a balance test that suggests conditional random assignment of intervention status to children in the sample. The $p$-value of the joint F-test is presented at the bottom, showing the (lack of) joint significance of the listed observable characteristics with respect to being assigned to the intervention group, conditioning on number of listed children for a subset of referrals with no mother and received after March 2021 (randomization procedure controls; Appendix C). Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
Table A.4: Robustness of Health Effects
Health Effect Change in Health Disparities
(1) (2)
@@
Chi-Squared Test p-value
All Four Disparities
1. Main Specification -0.061** <0.001<0.001
Permutation Test p-value [.04]
2SLS IV -0.083**
With Controls -0.056**
2. 60 Day Donut -0.067** <0.001<0.001
3. Including First 30 Days -0.051* <0.001<0.001
4. Full Sample -0.049* <0.001<0.001
5. O’Brien Weighted Index -0.054* <0.001<0.001
6. Index of Binary Variables -0.056 <0.001<0.001
7. First Principal Component -0.106** <0.001<0.001
8. Lower Bound - Treatment non-Compliance -0.047 <0.001<0.001
9. Upper Bound - Treatment non-Compliance -0.187*** <0.001<0.001
10. Clustering at the Referral Level -0.061** <0.001<0.001
Notes: This table presents the robustness of estimated reductions in the child harm index with algorithm support. Column 1 shows the main effect of the algorithm tool on children’s harm index, and column 2 shows a chi-squared test of whether harm disparities are unchanged by race, ethnicity, gender and socioeconomic status. Row 1 shows the benchmark estimates reported previously in the paper. Each subsequent row is a robustness test for these results. Below the first row, in brackets, we present a permutation test p-value where we randomly permute treatment status and account for randomization controls. We also report instrumental variables (IV) estimates of the effect of the algorithm, using our first-stage estimate of checking the algorithm score (recorded by workers; likely underestimate of true first stage). Finally, we report point estimates controlling for all variables listed in the balance test (Appendix Table A.3). In row 2, instead of excluding index hospitalizations during the pre-specified first 30 days after the referral, we exclude a larger window of 60 days within which most investigations have concluded. Row 3 includes all index hospitalizations after a referral, including within the first 30 days. In row 4, we report effects for the full sample instead of the hospital outcomes sample, meaning that we include children for whom the duration to examine hospital outcomes is extremely limited and predominantly falls after the trial. In row 5, we present a harm index that is not equally weighted on inputs, but rather uses weights that are optimal if the effect were to be the same on each of the outcome variables included in the index (O’Brien 1984). In row 6, instead of using the full discrete values of each of the five outcome variables, we construct an index that uses only the extensive margin (binary version) of each outcome. In row 7, instead of using the harm index as an outcome, we instead estimate the effects of algorithm access on the first principal component of a principal component analysis using the same five variables included in the harm index. In the final eighth and ninth rows, we present some approximate bounds to address treatment non-compliance (see Appendix C). In row 8, we replace the outcomes of children who ever changed treatment status to the minimum level of harm (no index hospital visits) for an extreme lower bound of treatment effect magnitudes. Estimates of access to the algorithm are still large and negative, suggesting that this small group of children do not drive the effects in our setting. Row 9 presents an extreme upper bound, showing that if these children all would have had the 99th percentile of the harm index, then estimates would be more than double the benchmark findings. Row 10 clusters standard errors at the referral level instead of at the household level. Unless otherwise specified, all regressions use the hospital outcomes sample, randomization procedure controls, and standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
Table A.5: Impacts of Algorithm Tool on Placebo Outcomes and Other Health Outcomes
Placebo Specifications Hospital Visits Medical Charges
(1) (2) (3) (4) (5) (6) (7) (8)
Harm when
Score Not Found
Cancer
Visits
Lower Priority
Visits
@@
N Suggestive
Maltreatment
(Schnitzer et al.)
Public Payer
Charges
Private Payer
Charges
Algorithm Available 0.023 0.0020 0.089 -0.020 -0.015 -0.0069 -42.7 -20.6
@@
Control Mean 0.000 0.003 1.242 0.471 1.902 0.019 925.747 116.853
Observations 1,787 3,431 3,431 3,431 3,431 1,769 3,431 3,431
Notes: This table reports the impact of algorithm access (i.e., estimates of equation 1) on the following: placebo specifications where the tool should have less of an impact (columns 1-3), any hospital visit, a child’s number of hospital visits, the number of ICD codes classified by Schnitzer et al. (2011) as suggestive of maltreatment for children aged 0-9 years old, public payer charges, and private payer charges. Charges are top-coded at the 95th percentile. In column 1 we report the effect on treated children’s harm index, compared to the control group, when workers tried to see the score but were unable (12 percent of treatment group). In column 2 we report the effect on cancer-related hospitalizations, a placebo outcome, which was pre-registered but extremely rare. Randomization procedure controls are included throughout. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
Table A.6: Effects of Algorithm Tool on CPS Outcomes
@@
Notes: This table reports the impact of algorithm access (i.e., estimates of equation 1) on whether a child was ever re-referred, the number of re-referrals, whether a child was ever removed to foster care, and the number of removal episodes. Estimates are presented for the full sample and are similar for the hospital outcomes sample (children randomized during the first year of the trial). Randomization procedure controls are included throughout. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
@@
Peer Share with Algorithm Available in Day 0.035 -0.126 0.039
@@
Algorithm Available $\cdot$ Peer Share -0.015 0.203 -0.046
@@
Mean of Outcome Variable 0.30 11.03 -0.04
Observations 4,127 3,267 3,067
Notes: This table tests for the presence of within-day spillovers (i.e., experimental contamination from treatment to control) from tool availability. If such within-day spillovers were present, one may expect changes in attention (screen-in or time) to children in the control group – and similar effects of a smaller magnitude for children in the treatment group – on days where workers saw a higher percent of children for whom the algorithm tool was available. The share of referrals with the tool available is constructed at the team-day level and excludes a child’s own referral (leave-one-out, i.e., jackknife share). Time to decision is constructed using timestamp data which is available for a large subset of referrals. Randomization procedure controls are included. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
Table A.8: Tool Availability Reduced Harm Disparities
@@
Algorithm Available $\cdot$ Type -0.23 -0.18 -0.061 -0.077
@@
Observations in Group 142 630 1,765 2,207
Observations 3,431 3,431 3,431 3,431
Effect on Disparity (%) -61 -84 -61 -53
p-value of Disparity Change 0.552 0.145 0.259 0.155
p-value of Chi-Squared Test <0.001<0.001
Notes: This table presents estimates of algorithm availability on disparities in the harm index for Black, Hispanic, female, and SNAP-recipient (lower income) children in columns 1-4, respectively, compared to other children in the sample. We report results from regressions of the child harm index on an interacted model of group type (reported in the column titles) with a binary indicator for whether the tool was available. For example, the coefficient in row 1 of column 1 indicates that Black children had a 0.37 greater harm index than other children when the tool was not available. Algorithm tool availability reduced harm by 0.047 standard deviations for other children (row 2) and an additional 0.23 for Black children (row 3). The sample is restricted to children in the hospital outcomes sample. The number of observations in each demographic group is listed at the bottom of the table. The table also reports the percentage decrease in the harm disparity based on point estimates. The second-to-last row from the very bottom lists the p-values from a test of whether health disparities are the same with and without the tool, and the final row reports the p-value of a joint test that none of the health disparities changed. Randomization procedure controls are included throughout. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
@@
Above Median Score (S) 0.048* 0.058* -0.071* 0.048* -0.077** 0.046 0.004*
(0.029) (0.031) (0.039) (0.029) (0.039) (0.029) (0.002)
@@
Randomization Controls Yes Yes Yes Yes Yes Yes Yes Yes
Reference Group Control Mean 0.267 0.267 0.261 0.418 0.267 0.418 0.268 0.294
Observations 4,681 4,681 3,865 3,321 4,681 4,681 4,681 4,681
Notes: This table shows estimates of racial disparities in screen-in across a variety of specifications. Columns 1 and 2 provide estimates on the change in disparities for Black vs. non-Black (or unknown race) screen-in rates when the algorithm tool is made available. We report permutation p-values in brackets in column 1. Column 3 is the same specification as column 1 but excludes Hispanic children. Column 4 excludes children where information about race was missing. Columns 5 and 6 recode children with missing race as white (column 5) or Black (column 6), respectively. Column 7 complements CPS-recorded race records with race information from hospital records. Finally, Column 8 replaces interaction with a binary indicator (above/below median algorithm score) with instead an interaction with a linear algorithmic score from 1-20 (pre-registered specification). Randomization procedure controls are included. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
@@
Workers mechanically follow tool predictions. No increase in CPS interventions for children with high algorithm risk scores (Figures A.11, A.12).
Workers exhibit algorithm aversion. No reduction in CPS intervention for children with high algorithm scores (Figures A.11, A.12). No qualitative evidence of tool aversion expressed by workers to researchers (independent from tool developers). Reductions in child harm (Table 2). Reductions in disparities (Figures 1, 2).
Experimenter demand effects. No differential screen-in (or other provision of services) by level of score (Figures A.11, A.12).
Workers reallocate CPS-adjacent services to high-risk children. Limited evidence of changes in levels/targeting of CPS-adjacent services (family visitor, prevention, law enforcement, community response) (Figures A.10, A.12). No evidence that these account for health effects.
Workers change behavior (e.g., exert more effort), for instance because they know they are being observed. Workers know they are being observed regardless of whether tool is shown. Tool access is random for each family and randomized within a day, so salience is comparable between intervention and control. No apparent spillovers (e.g., effort crowd-out) on control group outcomes on days with more treated families (Table A.7). Improvement in outcomes even for intermediate values of risk score (e.g., Figure A.6).
Investigator behavior (conditional on screen-in) influenced by tool. Tool not available to investigators unless they read through group meeting discussion notes. No evidence of changes in investigator behavior based on level of score. Improvement in outcomes for screened-out (non-investigated) children: Evidence that gains come at least in part from improved targeting of screen-ins prior to investigation (Table 3; Figures 3 and A.13).
@@
    Hospital Outcomes Sample. We had not foreseen that all families would become treated at the end of the trial, and were limited in the number of months of hospital data we could acquire after the trial. To minimize bias toward zero in treatment effects (which we had not considered during preregistration), we show results using hospital outcomes for those children randomized during the first year of the trial, which we call our “hospital outcomes sample” or “hospital sample” for short. We show in Table A.4 that results are consistent if we use the full sample.
@@
Each child has an unobserved maltreatment risk $r$, assumed in the context of this model to follow a normal distribution. [^18] The normal distribution provides closed-form solutions in contrast to alternative distributional assumptions. Assume for the purposes of this application that the mean and standard deviation of the normal distribution are such that $r$ is almost always positive. In the absence of a CPS intervention, a child’s realized harm is entirely determined by risk $r$. CPS workers can address this risk, however. Workers observe a noisy indicator of maltreatment risk, $m=r+\epsilon$, where $\epsilon$ is a normal error term with mean zero that is independent from $r$. Workers are constrained by state legal requirements in how they are instructed to respond to certain allegations (Appendix B2), so the model’s objective function is to minimize errors in estimating maltreatment risk based on the signal workers receive (i.e., accuracy in discerning risk) rather than solely minimizing child harm. Whereas underestimating risk could harm children, overestimating risk is costly in terms of worker time and effort, agency resources, and increases the risk of litigation. Hence, workers’ objective is to accurately assess child maltreatment risk.
@@
$=$ $\frac{1}{1-\Phi\left(\frac{-p^{*}}{\sigma_{p}}\right)}\left[\sigma_{p}\int_{\frac{-p^{*}}{\sigma_{p}}}^{+\infty}\phi^{\prime}(z)dz+p^{*}\int_{\frac{-p^{*}}{\sigma_{p}}}^{+\infty}\Phi^{\prime}(z)dz\right]$
$=$ $\frac{1}{1-\Phi\left(\frac{-p^{*}}{\sigma_{p}}\right)}\left[\sigma_{p}\left[\phi(z)\right]_{\frac{-p^{*}}{\sigma_{p}}}^{+\infty}+p^{*}\left[\Phi(z)\right]_{\frac{-p^{*}}{\sigma_{p}}}^{+\infty}\right]$
@@
Proof: First note that:
@@
$=\mathbb{P}(B)(1-\gamma^{T})\left[\mathbb{E}[p|T,B]-\alpha_{B}\right]=p^{*}<0$

### v3

@@ 2. Scored 13,115 7,021 7,345
3. First Scored Referral was in Trial 6,852 6,852 4,213
@@ 8. Final Sample, Excluding Observations with Overwritten Status 4,681 4,681 2,832
Table A.2: Sample Characteristics
Panel A: Child Demographics
Female 0.52
Age at Initial Randomization 10.5
White 0.66
Black 0.036
Missing Race 0.29
Hispanic 0.17
Missing Ethnicity 0.34
Panel B: Referral Details
N Children Listed on Referral 2.24
Neglect Allegation 0.77
Abuse Allegation 0.43
Panel C: Child Experiences
Total Referrals in Trial 2.19
Any Re-referral in Trial 0.42
Removed from Home during Trial 0.022
Any Hospital Visit After Referral 0.43
N Hospital Visits After Referral 1.68
N Unique Children 4,681
N Referrals 2,832
N Family Clusters 2,810
Notes: This table reports the mean characteristics of children who were reported to CPS and successfully randomized during the trial. Panels A and C are at the child level, while Panel B is at the referral level.
@@
Female 0.000
[0.014]
Age at Initial Randomization -0.002
[0.001]
White 0.048
[0.069]
Black -0.048
[0.082]
Missing Race -0.000
[0.077]
Hispanic 0.032
[0.029]
Missing Ethnicity 0.011
[0.037]
Algorithm Score -0.003
[0.002]
Mandated Reporter -0.011
[0.024]
Child Reported in Danger 0.038
[0.068]
Team 3 0.040
[0.025]
Team 2 0.005
[0.026]
Household Size 0.001
[0.014]
Number of Alleged Child Victims 0.000
[0.017]
Active -0.014
[0.034]
Number of Previous Referrals 0.001
[0.002]
Prior High Priority Hospital Visits 0.000
[0.018]
Prior Injury Hospital Visits 0.037
[0.030]
Prior Avoidable ER Hospital Visits 0.022
[0.036]
Prior Substance Exposure Hospital Visits 0.007
[0.032]
Prior Maltreatment ICD Hospital Visits 0.036
[0.135]
Prior Intentional Injury Hospital Visits -0.100
[0.070]
Any Prior Hospital Visit -0.038
[0.023]
Total Prior Hospital Visits 0.003
[0.004]
Referred in First Year of Trial 0.002
[0.023]
Randomization Controls Yes
Observations 4,681
Randomization Clusters 2,810
F-statistic 0.796
p-value of F-test 0.752
Notes: This table presents a balance test that suggests stratified random assignment of intervention status to children in the sample. The $p$-value of the joint F-test is presented at the bottom, showing the (lack of) joint significance of the listed observable characteristics with respect to being assigned to the intervention group, conditioning on number of listed children for a subset of referrals with no mother and received after March 2021 (randomization procedure controls; Appendix C). Referrals are assigned to one of three teams for processing. When a case or assessment is already open, the family is said to be “active.” All hospitalization variables are calculated on the 180 days prior to randomization to ensure the same time span prior to the first referral in the trial for all individuals. Apart from “Any Prior Hospital Visit” which is a binary measure, all other hospitalization variables are counts. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
Table A.5: Robustness of Harm Reduction Estimates
@@
Harm Index Injury
1. Main Specification -0.050** -0.034*
(-0.091, -0.009 ) (-0.065, -0.003 )
[p = 0.046] [p=0.068]
Permutation Test p-value [.041] [.068]
2SLS IV -0.069** -0.047*
(-0.125, -0.012 ) (-0.089, -0.005 )
[p=0.045] [p=0.068]
With Controls -0.046** -0.033*
(-0.082, -0.009 ) (-0.063, -0.003 )
[p=0.040] [p=0.068]
2. 60 Day Donut -0.055** -0.033*
(-0.096, -0.014 ) (-0.063, -0.004 )
[p = 0.027] [p=0.065]
3. Including First 30 Days -0.047* -0.041**
(-0.088, -0.007 ) (-0.074, -0.008 )
[p = 0.055] [p=0.042]
4. Count of Harm Visits -0.167*
(-0.307, -0.026 )
[p = 0.051]
5. Count of Distinct Harm Visits -0.074
(-0.149, 0.000 )
[p = 0.100]
6. O’Brien Weighted Index -0.047*
(-0.087, -0.008 )
[p = 0.050]
7. First Principal Component -0.094**
(-0.171, -0.017 )
[p = 0.045]
8. Lower Bound - Treatment non-Compliance -0.041 -0.029
(-0.082, 0.000 ) (-0.059, 0.002 )
[p = 0.101] [p=0.125]
9. Upper Bound - Treatment non-Compliance -0.140*** -0.102***
(-0.188, -0.092 ) (-0.138, -0.066 )
[p = 0.000] [p=0.000]
10. Clustering at the Referral Level -0.050** -0.034*
(-0.091, -0.009 ) (-0.065, -0.003 )
[p=0.045] [p=0.068]
Notes: This table presents the robustness of estimated reductions in the child harm index with algorithm support. Column 1 shows the main effect of the algorithm tool on children’s harm index, and column 2 shows effects for injury-related hospitalizations. Row 1 presents benchmark estimates reported previously in the paper. Each subsequent row is a robustness test for these results. Below the first row, in brackets, we present a permutation test p-value where we randomly permute treatment status and account for randomization controls. We also report instrumental variables (IV) estimates of the effect of the algorithm, using our first-stage estimate of checking the algorithm score (recorded by workers; likely underestimate of true first stage). Finally, we report point estimates controlling for all variables, including prior hospitalizations, listed in the balance test (Appendix Table A.4). In row 2, instead of excluding index hospitalizations during the pre-specified first 30 days after the referral, we exclude a larger window of 60 days within which most investigations have concluded. Row 3 includes all index hospitalizations after a referral, including within the first 30 days. In row 4, we report effects on the sum of all harm-index-related hospitalizations, where a hospitalization for multiple categories is counted as multiple visits. In row 5, we report effects on a count of harm-index-related hospitalizations where overlapping categories are only counted once (e.g., a hospitalization for both substance exposure and injuries is counted as a single visit). In practice, this outcome is almost identical to our most frequent outcome, the number of high priority visits (see text). In row 6, we present a harm index that is not equally weighted on inputs, but rather uses weights that are optimal if the effect were to be the same on each of the outcome variables included in the index (O’Brien 1984). In row 7, instead of using the harm index as an outcome, we instead estimate the effects of algorithm access on the first principal component of a principal component analysis using the same six variables included in the harm index. In the final eighth and ninth rows, we present some approximate bounds to address treatment non-compliance (see Appendix C). In row 8, we replace the outcomes of children who ever changed treatment status to the minimum level of harm (no index hospital visits) for an extreme lower bound of treatment effect magnitudes. Estimates of access to the algorithm are still large and negative, suggesting that this small group of children do not drive the effects in our setting. Row 9 presents an extreme upper bound, showing that if these children all would have had the 99th percentile of the harm index, then estimates would be more than double the benchmark findings. The final row clusters standard errors at the referral level instead of at the household level. All regressions include randomization procedure controls, and standard errors are clustered at the household level unless otherwise specified. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
Table A.6: Effects on Placebo Outcomes and Other Health Outcomes
Hospital Visits Placebo Specifications
(1) (2) (3) (4) (5) (6) (7) (8)
N Suggestive
Maltreatment
(Schnitzer et al.)
@@
Extensive Margin
@@
Public+Self
Payer Charges
Cancer
Visits
Harm when
Score Not Found
Lower Priority
Visits
Algorithm Available -0.0088 -0.038 -0.0053 -0.023 -199.8 0.0012 0.020 0.046
@@
Control Mean 0.018 -0.000 0.412 1.586 1,809.463 0.003 -0.000 1.042
Effect (%) -50 -1 -1 -11 44 4
Design Power 0.69 0.84 0.63 0.26 0.63 0.09 0.13 0.09
Observations 2,445 4,681 4,681 4,681 4,681 4,681 2,393 4,681
Notes: This table reports the impact of algorithm access (i.e., estimates of equation 1) on the following: the number of ICD codes classified by Schnitzer et al. (2011) as suggestive of maltreatment for children aged 0-9 years old; an extensive margin harm index where rather than using the number of visits for each component of the harm index we instead use binary indicators for whether there was any such visit after the referral; any hospital visit; a child’s number of hospital visits; and public or self payer charges (columns 1-5). Charges are top-coded at the 95th percentile. Placebo specifications where the tool should have no impact are reported in columns 6-8. In column 6 we report the effect on cancer-related hospitalizations, a placebo outcome which was pre-registered but extremely rare. In column 7 we report the effect on treated children’s harm index relative to the control group, when workers tried to see the score but were unable (11 percent of treatment group). Finally, column 8 reports impacts on lower-priority hospital visits, which are less likely to be related to maltreatment and should be less affected by the tool. Randomization procedure controls are included throughout. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
Table A.7: Effect of Algorithm Tool on CPS Outcomes
@@
Notes: This table reports the impact of algorithm access (i.e., estimates of equation 1) on whether a child was ever re-referred, the number of re-referrals, whether a child was ever removed to foster care, and the number of removal episodes. Randomization procedure controls are included throughout. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
@@
Peer Share with Algorithm Available in Day (Std Dev) -0.005 -0.387 0.001
@@
Algorithm Available $\cdot$ Peer Share (Std Dev) 0.021 0.427 -0.018
@@
Mean of Outcome Variable 0.30 20.01 -0.04
Observations 4,501 4,501 4,501
Notes: This table tests for the presence of within-day spillovers (i.e., experimental contamination from treatment to control) from tool availability. If such within-day spillovers were present, one may expect changes in attention (screen-in or time) to children in the control group – and similar effects of a smaller magnitude for children in the treatment group – on days where workers saw a higher percent of children for whom the algorithm tool was available. The share of referrals with the tool available is constructed at the team-day level and excludes a child’s own referral (leave-one-out, i.e., jackknife share). Days with only one referral are excluded. Time to decision is constructed using timestamp data which is available for a large subset of referrals and is top-coded at 60 minutes. Randomization procedure controls are included. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
Table A.9: Effect on Harm Disparities
@@
Algorithm Available $\cdot$ Type -0.23 -0.15 -0.074 -0.061
@@
Observations in Group 170 816 2,452 2,977
Observations 4,681 4,681 4,681 4,681
Effect on Disparity (%) -62 -82 -80 -55
p-value of Disparity Change 0.546 0.164 0.101 0.172
p-value of Joint Test .2136
Notes: This table presents estimates of algorithm availability on disparities in the harm index for Black, Hispanic, female, and SNAP-recipient (lower income) children in columns 1-4, respectively, compared to other children in the sample. We report results from regressions of the child harm index on an interacted model of group type (reported in the column titles) with a binary indicator for whether the tool was available. For example, the coefficient in row 1 of column 1 indicates that Black children had a 0.38 greater harm index than other children when the tool was not available. Algorithm tool availability reduced harm by 0.038 standard deviations for other children (row 2) and an additional 0.23 for Black children (row 3). The number of observations in each demographic group is listed at the bottom of the table. The table also reports the percentage decrease in the harm disparity based on point estimates. The second-to-last row from the very bottom lists the p-values from a test of whether health disparities are the same with and without the tool, and the final row reports the p-value of a joint test that none of the health disparities changed. Randomization procedure controls are included throughout. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
@@
Above Median Score (S) 0.048* 0.058* -0.071* -0.077** 0.046 0.004*
(0.029) (0.031) (0.039) (0.039) (0.029) (0.002)
@@
Randomization Controls Yes Yes Yes Yes Yes Yes Yes
Reference Group Control Mean 0.267 0.267 0.261 0.418 0.418 0.268 0.294
Observations 4,681 4,681 3,865 3,321 4,681 4,681 4,681
Notes: This table shows estimates of racial disparities in screen-in across a variety of specifications. Columns 1 and 2 provide estimates on the change in disparities for Black vs. non-Black (or unknown race) screen-in rates when the algorithm tool is made available. Columns 1 and 2 code missing race as white. We report permutation p-values in brackets in column 1. Column 3 is the same specification as column 1 but excludes Hispanic children. Column 4 excludes children where information about race was missing. Column 5 recodes children with missing race as Black. Column 6 complements CPS-recorded race records with race information from hospital records. Finally, Column 7 replaces interaction with a binary indicator (above/below median algorithm score) with instead an interaction with a linear algorithm score from 1-20 (pre-registered specification). Randomization procedure controls are included. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
@@
Workers mechanically follow tool predictions. No increase in CPS interventions for children with high algorithm risk scores (Figures A.12, A.13).
Workers exhibit algorithm aversion. No reduction in CPS intervention for children with high algorithm scores (Figures A.12, A.13). No qualitative evidence of tool aversion expressed by workers to researchers (independent from tool developers). Reductions in child harm (Table 1). Reductions in disparities (Figures 1, 2).
Experimenter demand effects. No differential screen-in (or other provision of services) by level of score (Figures A.12, A.13).
Workers reallocate CPS-adjacent services to high-risk children. Limited evidence of changes in levels/targeting of CPS-adjacent services (family visitor, prevention, law enforcement, community response) (Figures A.11, A.13). No evidence that these account for health effects.
Workers change behavior (e.g., exert more effort), for instance because they know they are being observed. Workers know they are being observed regardless of whether tool is shown. Tool access is random for each family and randomized within a day, so salience is comparable between intervention and control. No apparent spillovers (e.g., effort crowd-out) on control group outcomes on days with more treated families (Table A.8).
Investigator behavior (conditional on screen-in) influenced by tool. Tool not available to investigators unless they read through group meeting discussion notes. No evidence of changes in investigator behavior based on level of score. Improvement in outcomes for screened-out (non-investigated) children: Evidence that gains come at least in part from improved targeting of screen-ins prior to investigation (Table 2; Figures 3 and A.14).
@@
    Suggestive Maltreatment Outcome. Among ICD codes suggestive of maltreatment, we had included a vague reference to the trauma mortality prediction model (TMPM). However, the inconsistent level of ICD-code detail in our received data, combined with difficulty of crosswalks between ICD-9 to ICD-10, made it impossible to apply the TMPM to our setting. We thus report effects on an alternative measure of suggestive maltreatment provided by Schnitzer et al. (2011) in Appendix Table A.6. As would have been the case for the TMPM, this outcome is only available for a subset of child ages.
@@
Each child has an unobserved maltreatment risk $r$, assumed in the context of this model to follow a normal distribution. [^23] The normal distribution provides closed-form solutions in contrast to alternative distributional assumptions. Assume for the purposes of this application that the mean and standard deviation of the normal distribution are such that $r$ is almost always positive. In the absence of a CPS intervention, a child’s realized harm is entirely determined by risk $r$. CPS workers can address this risk, however.
Workers observe a noisy indicator of maltreatment risk, $m=r+\epsilon$, where $\epsilon$ is a normal error term with mean zero that is independent from $r$. Workers are constrained by state legal requirements in how they are instructed to respond to certain allegations (Appendix B2), so the model’s objective function is to minimize errors in estimating maltreatment risk based on the signal workers receive (i.e., accuracy in discerning risk) rather than solely minimizing child harm. Whereas underestimating risk could harm children, overestimating risk is costly in terms of worker time and effort, agency resources, and increases the risk of litigation. Hence, workers’ objective is to accurately assess child maltreatment risk.
@@
$=$ $\frac{1}{1-\Phi\left(\frac{-p^{*}}{\sigma_{p}}\right)}\left[\sigma_{p}\int_{\frac{-p^{*}}{\sigma_{p}}}^{+\infty}-\phi^{\prime}(z)dz+p^{*}\int_{\frac{-p^{*}}{\sigma_{p}}}^{+\infty}\Phi^{\prime}(z)dz\right]$
$=$ $\frac{1}{1-\Phi\left(\frac{-p^{*}}{\sigma_{p}}\right)}\left[\sigma_{p}\left[-\phi(z)\right]_{\frac{-p^{*}}{\sigma_{p}}}^{+\infty}+p^{*}\left[\Phi(z)\right]_{\frac{-p^{*}}{\sigma_{p}}}^{+\infty}\right]$
@@
Proof: First note that, since predictions are unbiased for group W and then replacing $p$ by its expression, we get:
@@
$=\mathbb{P}(B)(1-\gamma^{T})\left[\mathbb{E}[r|T,B]-\alpha_{B}\right]$
$=\mathbb{P}(B)(1-\gamma^{T})\left[\mathbb{E}[r|B]-\alpha_{B}\right]=p^{*}<0$